OpenStudy (anonymous):

\[(x ^{3/5})^{2/3}\] \[(x ^{3/5})^{2/3}\] @Mathematics

OpenStudy (anonymous):

multiply the powers..

OpenStudy (anonymous):

^6/15?

OpenStudy (anonymous):

yes u can simplify it 6/15=2/5

OpenStudy (anonymous):

how would i do this...

OpenStudy (anonymous):

\[(x ^{3})^{3/4}\]

OpenStudy (hoblos):

x^9/3

OpenStudy (anonymous):

\[x^{6/15}=x ^{2/5}\]

OpenStudy (anonymous):

i have to put it in radical form

OpenStudy (anonymous):

which one now?

OpenStudy (anonymous):

the second problem.

OpenStudy (anonymous):

(x^3)^3/4

OpenStudy (anonymous):

\[x ^{9/4}=\sqrt[4]{x ^{9}}=\sqrt[4]{x ^{5}x ^{4}}=x \sqrt[4]{x ^{5}}\]

OpenStudy (anonymous):

but thats not one of my options.

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

\[x \sqrt[4]{x ^{5}}=x^{2}\sqrt[4]{x}\]

OpenStudy (anonymous):

this one??